# Hui Xiong2, Bin Yang1

### v1

@@ 3.3 Time-Frequency Training
where $\tau$ is the temperature coefficient. The ClusteringLoss is similar to the InfoNCE (He et al., 2020) in form but does not fix the number of “positive” pairs. In contrast, it changes its “positive” pairs for different patches based on the current discovered channel correlation $\mathcal{M}^{i}$. As shown in Figure 2, it sets the Query and Key views of relevent channels in a frequency band ($\mathcal{M}^{i}_{l,m}=1$) as the “positive” pairs, thus encouraging the $\mathrm{W}^{Q}$ and $\mathrm{W}^{K}$ to cluster the patch-wise relevent channels in the hidden spaces and lead to higher attention scores. We share the calculating results of $\mathcal{S}^{i}$ and $\mathcal{T}^{i}$ with the attention mechanism in Section 3.2 to save the computational cost. However, only a single ClusteringLoss may cause some adverse effects by urging the mask generator to output constant ones matrix, so that we add a RegularLoss to mitigate this risk by restricting the number of relevent channels. Equipped with the two optimization objectives, the mask generator is encouraged to discover appropriate patch-wise channel correlations between CI and CD, and the attention mechanism is also enchanced by optimizing the $\mathrm{W}^{Q}$ and $\mathrm{W}^{K}$ to learn fine-grained channel representations in the hidden spaces.
@@
Time-Frequency Reconstruction. We adopt the reconstruction loss functions both in time and frequency domains to separately enhance the ability of point-to-point and subsequence modeling. The loss functions for input series $\mathrm{X}$ is formalized as:
$\text{RecLoss}^{time}=||\mathrm{X}-\mathrm{X^{\prime}}||_{F}^{2},\text{RecLoss}^{freq}=||\mathrm{X}^{R}-\mathrm{X^{\prime}}^{R}||_{1}+||\mathrm{X}^{I}-\mathrm{X^{\prime}}^{I}||_{1}$ (11)
@@
We utilize $2$-norm in the time domain and $1$-norm in the frequency domain due to the distinct numerical characteristics of time and frequency domains (Wang et al., 2024).
Bi-level optimization. We weightsum the four optimization objectives to a TotalLoss $\mathcal{L}$ and utilize a bi-level optimization Algorithm 1 to iteratively update the mask generator and other model parameters. Intuitively, the process optimizes model parameters for current channel correlations and then discovers better channel correlations for the optimized model parameters, which facilitates the refinement of channel correlations in a continuous way.
@@
1:Input: Model parameters $\theta_{\text{model}}$, $\theta_{\text{mask}}$, learning rate $\eta_{\text{model}}$, $\eta_{\text{mask}}$, number of iterations $\mathcal{N}_{O}$, $\mathcal{N}_{I}$, loss function $\mathcal{L}$ = $\lambda_{1}\cdot\text{RecLoss}^{time}$ +$\lambda_{2}\cdot\text{RecLoss}^{freq}$ + $\lambda_{3}\cdot\text{ClusteringLoss}$ + $\lambda_{4}\cdot\text{RegularLoss}$
@@
3.4 Frequency-Enhanced Point-Granularity Scoring
When calculating the anomaly score, the convention is to obey the point-to-point manner by assigning an anomaly score for each timestamp, thus mainly reflecting the point anomlies in the time domain. To better quantify subsequence anomalies, existing methods often consider coarse-grained window-granularity scoring, which adds a frequency anomaly score to each point in the whole input window. However, they fail to know the actual boundaries of subsequence anomalies, thus causing misjudgment or omission. As our proposed CATCH is optimized for $\text{RecLoss}^{freq}$, it is capable of detecting subsequence anomalies in the frequency domain. During scoring, we perform the patching operation in the input window with the stride length equal to $1$.
Figure 3: Proposed Anomaly Scoring.
In Figure 3, the shadow in the Time Series indicates a series of subsequence anomalies. Take the red point as an example, we collect all the patches to which it belongs, then transform these patches into the frequency domain, calculate the reconstructed frequency anomaly scores through $\text{RecLoss}^{freq}$ and take the average value as frequency anomaly score of this point, as this gives superior performance compared to the minimum or the maximum (Schmidl et al., 2022). Obviously, the method can reflect the real surroundings of each point by considering all possible subsequence anomalies around this point, thus achieving the point-granularity alignment and showing strong sensitivity (Nam et al., 2024). Finally, we weightsum the time and frequency anomaly scores point-to-point. We provide the details of an efficient implementation version in Appendix A.3.
@@ 4 Experiments
Datasets We conduct experiments using 9 real-world datasets and 12 synthetic datasets (TODS datasets) to assess the performance of CATCH, more details of the benchmark datasets are included in Appendix A.1. The synthetic datasets are generated using the method reported in (Lai et al., 2021). Please refer to Appendix A.5 for specific implementation. We report the results on 9 real-world MTSAD datasets, including MSL, PSM, SMD, CICIDS, CalIt2, NYC, Creditcard, GECCO and Genesis in the main text. We also report the mean results of the 6 types of synthetic anomalies. The complete results can be found in the Appendix C.
Baselines We comprehensively compare our model against 14 baselines, including the latest state-of-the-art (SOTA) models. These baselines feature the 2024 SOTA iTransformer (iTrans) (Liu et al., 2023), ModernTCN (Modern) (Luo & Wang, 2024), and DualTF (Nam et al., 2024), along with the 2023 SOTA Anomaly Transformer (ATrans) (Xu et al., 2021), DCdetector (DC) (Yang et al., 2023b), TimesNet (TsNet) (Wu et al., 2022), PatchTST (Patch) (Nie et al., 2022), DLinear (DLin) (Zeng et al., 2023), NLinear (NLin) (Zeng et al., 2023), and AutoEncoder (AE) Sakurada & Yairi (2014). Additionally, we include non-learning methods such as One-Class SVM (OCSVM)(Schölkopf et al., 1999), Isolation Forest (IF)(Liu et al., 2008), Principal Component Analysis (PCA) (Shyu et al., 2003), and HBOS (Goldstein & Dengel, 2012).
@@
Table 1: Average A-R (AUC-ROC) and Aff-F (Affiliated-F1) accuracy measures for 9 real-world datasets and 6 synthetic datasets of different types of anomalies. The best results are highlighted in bold, and the second-best results are underlined.
Dataset Metric CATCH Modern iTrans DualTF ATrans DC TsNet Patch DLin NLin AE Ocsvm IF PCA HBOS
CICIDS Aff-F 0.787 0.654 0.708 0.692 0.560 0.664 0.657 0.660 0.669 0.669 0.243 0.693 0.604 0.619 0.542
A-R 0.795 0.697 0.692 0.603 0.528 0.638 0.732 0.716 0.751 0.691 0.629 0.537 0.787 0.601 0.760
CalIt2 Aff-F 0.835 0.780 0.812 0.751 0.729 0.697 0.794 0.793 0.793 0.757 0.587 0.783 0.402 0.768 0.756
A-R 0.838 0.676 0.791 0.574 0.533 0.527 0.771 0.808 0.752 0.695 0.767 0.804 0.775 0.790 0.798
Credit Aff-F 0.750 0.744 0.713 0.663 0.650 0.632 0.744 0.746 0.738 0.742 0.561 0.714 0.634 0.710 0.695
A-R 0.958 0.957 0.934 0.703 0.552 0.504 0.957 0.957 0.954 0.948 0.909 0.953 0.860 0.871 0.951
GECCO Aff-F 0.908 0.893 0.839 0.701 0.782 0.687 0.894 0.906 0.893 0.882 0.823 0.666 0.424 0.785 0.708
A-R 0.970 0.952 0.795 0.714 0.516 0.555 0.954 0.949 0.947 0.936 0.769 0.804 0.619 0.711 0.557
Genesis Aff-F 0.896 0.833 0.891 0.810 0.856 0.776 0.864 0.856 0.856 0.829 0.854 0.677 0.788 0.814 0.721
A-R 0.974 0.676 0.690 0.937 0.947 0.659 0.913 0.685 0.696 0.755 0.931 0.733 0.549 0.815 0.897
MSL Aff-F 0.740 0.726 0.710 0.588 0.692 0.694 0.734 0.724 0.725 0.723 0.625 0.641 0.584 0.678 0.680
A-R 0.664 0.633 0.611 0.576 0.508 0.507 0.613 0.637 0.624 0.592 0.562 0.524 0.524 0.552 0.574
NYC Aff-F 0.994 0.769 0.684 0.708 0.853 0.862 0.794 0.776 0.828 0.819 0.689 0.667 0.648 0.680 0.675
A-R 0.816 0.466 0.640 0.633 0.671 0.549 0.791 0.709 0.768 0.671 0.504 0.456 0.475 0.666 0.446
PSM Aff-F 0.859 0.825 0.854 0.725 0.710 0.682 0.842 0.831 0.831 0.843 0.707 0.531 0.620 0.702 0.658
A-R 0.652 0.593 0.592 0.600 0.514 0.501 0.592 0.586 0.580 0.585 0.650 0.619 0.542 0.648 0.620
SMD Aff-F 0.847 0.840 0.827 0.679 0.724 0.675 0.831 0.845 0.841 0.844 0.439 0.742 0.626 0.738 0.629
A-R 0.811 0.722 0.745 0.631 0.508 0.502 0.727 0.736 0.728 0.738 0.774 0.602 0.664 0.679 0.626
Contextual Aff-F 0.823 0.619 0.802 0.635 0.601 0.597 0.666 0.766 0.780 0.700 0.755 0.696 0.679 0.475 0.481
A-R 0.910 0.562 0.905 0.598 0.546 0.525 0.908 0.854 0.700 0.530 0.896 0.711 0.821 0.538 0.464
Global Aff-F 0.949 0.748 0.922 0.649 0.656 0.567 0.910 0.940 0.928 0.808 0.919 0.849 0.912 0.704 0.528
A-R 0.997 0.873 0.976 0.595 0.564 0.514 0.989 0.992 0.979 0.675 0.996 0.996 0.938 0.758 0.608
Seasonal Aff-F 0.997 0.681 0.992 0.776 0.788 0.859 0.992 0.989 0.993 0.951 0.927 0.805 0.938 0.637 0.673
A-R 0.998 0.512 0.946 0.701 0.584 0.644 0.958 0.922 0.823 0.623 0.949 0.829 0.918 0.437 0.516
Shapelet Aff-F 0.985 0.675 0.961 0.692 0.699 0.737 0.941 0.933 0.961 0.759 0.871 0.771 0.887 0.683 0.640
A-R 0.970 0.522 0.864 0.573 0.519 0.597 0.877 0.818 0.684 0.563 0.865 0.655 0.748 0.517 0.337
Trend Aff-F 0.916 0.734 0.901 0.677 0.584 0.765 0.897 0.888 0.721 0.830 0.699 0.691 0.914 0.693 0.669
A-R 0.892 0.612 0.847 0.524 0.500 0.569 0.858 0.835 0.671 0.642 0.482 0.471 0.878 0.484 0.468
Mixture Aff-F 0.892 0.856 0.862 0.652 0.641 0.709 0.863 0.879 0.727 0.839 0.673 0.676 0.881 0.676 0.667
A-R 0.931 0.763 0.854 0.570 0.522 0.516 0.861 0.863 0.767 0.749 0.493 0.475 0.911 0.517 0.531
Table 2: Multi-metrics results on three real-world multivariate datasets. The best ones are in Bold.
Dataset Method Acc P R F1 R-P R-R R-F Aff-P Aff-R Aff-F A-R A-P R-A-R R-A-P V-ROC V-PR
GECCO TimesNet 0.984 0.379 0.804 0.516 0.053 0.782 0.099 0.810 0.997 0.894 0.954 0.410 0.977 0.428 0.974 0.429
ModernTCN 0.984 0.373 0.779 0.504 0.086 0.644 0.152 0.808 0.998 0.893 0.952 0.447 0.978 0.459 0.975 0.461
CATCH (ours) 0.984 0.380 0.818 0.518 0.065 0.795 0.119 0.832 0.998 0.908 0.970 0.418 0.990 0.473 0.987 0.465
MSL TimesNet 0.855 0.166 0.093 0.119 0.130 0.224 0.164 0.589 0.973 0.734 0.613 0.146 0.701 0.231 0.692 0.227
ModernTCN 0.857 0.166 0.090 0.117 0.129 0.194 0.155 0.578 0.975 0.726 0.633 0.146 0.708 0.224 0.701 0.220
CATCH (ours) 0.853 0.185 0.117 0.143 0.150 0.241 0.185 0.599 0.966 0.740 0.664 0.167 0.747 0.260 0.735 0.256
SMD TimesNet 0.931 0.176 0.181 0.178 0.110 0.385 0.171 0.745 0.938 0.831 0.727 0.141 0.747 0.140 0.746 0.140
ModernTCN 0.931 0.151 0.145 0.148 0.092 0.378 0.148 0.755 0.948 0.840 0.722 0.130 0.743 0.130 0.742 0.130
CATCH (ours) 0.918 0.194 0.305 0.237 0.095 0.478 0.158 0.773 0.938 0.847 0.811 0.172 0.800 0.159 0.797 0.159
@@ 4.1 Main Results
We first evaluate CATCH with 14 competitive baselines on 9 real-world multivariate and 6 types of synthetic multivariate datasets generated by the methods reported in TODS (Lai et al., 2021) as shown in Table 1. It can be seen that our proposed CATCH achieves SOTA results under the widely used Affiliated-F1-score metric in all benchmark datasets. Besides, CATCH has the highest AUC-ROC values on all datasets. It means that our model performs well in the false-positive and true-positive rates under various pre-selected thresholds, which is important for real-world applications. CATCH effectively handles both point (Contextual, Global) and subsequence (Seasonal, Shapelet, Trend, Mixture) anomalies while showing greater improvement in detecting the subsequence anomalies. In addition, as shown in Table 2, CATCH’s performance on other metrics mostly remains at the forefront, further validating the robustness of our algorithm.

### v3

@@ 3.2 Channel Fusion Module
The ClusteringLoss is similar to the InfoNCE (He et al., 2020) in form but does not fix the number of “positive” pairs. In contrast, it changes its “positive” pairs for different patches based on the current discovered channel correlation $\mathcal{M}^{i}$. Theoretically, the similarities between “positive” or “negative” pairs are also calculated through inner product, so we share the calculating results of $\mathcal{S}^{i}$ and $\mathcal{T}^{i}$ from the attention mechanism to save the computational cost. As shown in Figure 2, we exhibit the “positive” and “negative” pairs in the attention matrix $\mathcal{T}$ and masked attention matrix $\mathcal{S}$. Specifically, it sets the Query and Key views of relevant channels in a frequency band ($\mathcal{M}^{i}_{l,m}=1$) as the “positive” pairs, otherwise ($\mathcal{M}^{i}_{l,m}=0$) “negative” pairs, thus encouraging the $\mathrm{W}^{Q}$ and $\mathrm{W}^{K}$ to cluster the patch-wise relevant channels in the hidden spaces and lead to higher attention scores. However, only a single ClusteringLoss may cause some adverse effects by urging the Mask Generator to output constant ones matrix, so that we add a RegularLoss to mitigate this risk by restricting the number of relevant channels. Equipped with the two optimization objectives, the Mask Generator is encouraged to discover appropriate patch-wise channel correlations between CI and CD, and the attention mechanism is also enchanced by optimizing the $\mathrm{W}^{Q}$ and $\mathrm{W}^{K}$ to learn fine-grained channel representations in the hidden spaces.
3.3 Time Frequency Reconstruction Module
The Time-Frequency Reconstruction Module (TFRM) contains the following two components: 1) the Flatten & Linear Head Layer and 2) the iFFT Layer. Specifically, after the CFM fully extract the fine-grained channel correlations, we utilize the TFRM to flatten the patch-wise representations and reconstruct all frequency spectrums with MLP projections separately for real and imaginary patches, and obtain temporal reconstruction through iFFT:
$\mathrm{X^{\prime}}$ $=\text{iFFT}(\mathrm{X^{\prime}}^{R},\mathrm{X^{\prime}}^{I}),$ (13)
@@ 3.3 Time Frequency Reconstruction Module
where $\mathrm{X^{\prime}}^{R}$, $\mathrm{X^{\prime}}^{I}$, and $\mathrm{X^{\prime}}\in\mathbb{R}^{N\times T}$. We then adopt the reconstruction loss functions both in time and frequency domains to separately enhance the ability of point-to-point and subsequence modeling. The reconstruction functions in time and frequency domains are formalized as: $\text{RecLoss}^{time}=||\mathrm{X}-\mathrm{X^{\prime}}||_{F}^{2}$ (14) $\text{RecLoss}^{freq}=||\mathrm{X}^{R}-\mathrm{X^{\prime}}^{R}||_{1}+||\mathrm{X}^{I}-\mathrm{X^{\prime}}^{I}||_{1}$ (15)
@@
We utilize $2$-norm in the time domain and $1$-norm in the frequency domain due to the distinct numerical characteristics of time and frequency domains (Wang et al., 2024d).
3.4 Joint Bi-level optimization
We then design a novel joint bi-level training process to enhance the model’s ability to detect both point anomalies and subsequence anomalies. Our TotalLoss $\mathcal{L}$ mainly consists of reconstruction loss functions in the time ($\text{RecLoss}^{time}$) and frequency ($\text{RecLoss}^{freq}$) domains, ClusteringLoss and RegularLoss from the Channel Correlation Discovering mechanism. The reconstruction loss functions are used to enhance model’s ability in both time and frequency domains to detect point and subsequence anomlies. The ClusteringLoss and RegularLoss are used to guide the discovering for fine-grained channel correlations. Specifically, we weightsum these four optimization objectives:
@@ 3.4 Joint Bi-level optimization
1:Input: Model parameters $\theta_{\text{model}}$, $\theta_{\text{mask}}$, learning rate $\eta_{\text{model}}$, $\eta_{\text{mask}}$, number of iterations $\mathcal{N}_{O}$, $\mathcal{N}_{I}$, loss function $\mathcal{L}$ = $\text{RecLoss}^{time}$ + $\lambda_{1}\cdot\text{RecLoss}^{freq}$ + $\lambda_{2}\cdot\text{ClusteringLoss}$ + $\lambda_{3}\cdot\text{RegularLoss}$
@@
where $\lambda_{1},\lambda_{2},$ and $\lambda_{3}$ are empircal coefficients. We then utilize a bi-level optimization Algorithm 1 to iteratively update the Mask Generator and other model parameters. Intuitively, the process optimizes model parameters for current channel correlations and then discovers better channel correlations for the optimized model parameters, which facilitates the refinement of channel correlations in a continuous way.
3.5 Anomaly scoring
When calculating the anomaly score, the convention is to obey the point-to-point manner by assigning an anomaly score for each timestamp, thus mainly reflecting the point anomlies in the time domain. To better quantify subsequence anomalies,
Figure 3: Anomaly Scoring.
existing methods often consider coarse-grained window-granularity scoring, which adds a frequency anomaly score to each point in the whole input window (Ren et al., 2019; Park et al., 2021; Zhang et al., 2022). However, they fail to know the actual boundaries of subsequence anomalies, thus causing misjudgment or omission. During scoring, we perform the patching operation in the input window with the stride length equal to $1$. In Figure 3, the shadow in the Time Series indicates a series of subsequence anomalies. Take the calculation of the anomaly score for the red time point in the Time Series as an example, we first calculate the time domain reconstruction error ($\text{RecLoss}^{time}$) between its reconstructed value in the time domain and the true value using Equation 14, and denote the result as the time domain anomaly score (time-score). At the same time, we collect all patches containing the red time point, transform these patches into the frequency domain, and compute the frequency reconstruction error ($\text{RecLoss}^{freq}$) between the reconstructed value and the true frequency-domain value using Equation 15. The average frequency reconstruction error of all patches is then taken as the frequency domain anomaly score (freq-score) for this red time point, as the average performs better than the minimum or maximum value (Schmidl et al., 2022). Finally, we weightsum the time-score and freq-score point-to-point to obtain the final anomaly score for the Time Series.
[…]
For ease of understanding, we provide an efficient implementation version of the calculation of freq-score in Appendix A.5. Obviously, our method can reflect the real surroundings of each point by considering all possible subsequence anomalies around this point, thus achieving the point-granularity alignment and showing strong sensitivity (Nam et al., 2024).
@@ 4 Experiments
Datasets We conduct experiments using 10 real-world datasets and 12 synthetic datasets (TODS datasets) to assess the performance of CATCH, more details of the benchmark datasets are included in Appendix A.1. The synthetic datasets are generated using the method reported in (Lai et al., 2021). Please refer to Appendix A.6 for specific implementation. We report the results on 10 real-world MTSAD datasets, including MSL, PSM, SMD, CICIDS, CalIt2, NYC, Creditcard, GECCO, Genesis, and ASD in the main text. We also report the mean results of the 6 types of synthetic anomalies. The complete results can be found in the Appendix C.
Baselines We comprehensively compare our model against 15 baselines, including the latest state-of-the-art (SOTA) models. These baselines feature the 2024 SOTA iTransformer (iTrans) (Liu et al., 2024c), ModernTCN (Modern) (Luo & Wang, 2024), and DualTF (Nam et al., 2024), along with the 2023 SOTA Anomaly Transformer (ATrans) (Xu et al., 2021), DCdetector (DC) (Yang et al., 2023b), TimesNet (TsNet) (Wu et al., 2023), PatchTST (Patch) (Nie et al., 2023), DLinear (DLin) (Zeng et al., 2023), NLinear (NLin) (Zeng et al., 2023), TFAD (Zhang et al., 2022), and AutoEncoder (AE) (Sakurada & Yairi, 2014). Additionally, we include non-learning methods such as One-Class SVM (OCSVM)(Schölkopf et al., 1999), Isolation Forest (IF)(Liu et al., 2008), Principal Component Analysis (PCA) (Shyu et al., 2003), and HBOS (Goldstein & Dengel, 2012).
@@ 4.1 Main Results
We first evaluate CATCH with 15 competitive baselines on 10 real-world multivariate and 6 types of synthetic multivariate datasets generated by the methods reported in TODS (Lai et al., 2021) as shown in Table 2. It can be seen that our proposed CATCH achieves SOTA results under the widely used Affiliated-F1-score metric in all benchmark datasets. Besides, CATCH has the highest AUC-ROC values on all datasets. It means that our model performs well in the false-positive and true-positive rates under various pre-selected thresholds, which is important for real-world applications. CATCH effectively handles both point (Contextual, Global) and subsequence (Seasonal, Shapelet, Trend, Mixture) anomalies while showing greater improvement in detecting the subsequence anomalies.
Table 2: Average A-R (AUC-ROC) and Aff-F (Affiliated-F1) accuracy measures for 10 real-world datasets and 6 synthetic datasets of different types of anomalies. The best results are highlighted in bold, and the second-best results are underlined.
Dataset Metric CATCH Modern iTrans DualTF ATrans DC TsNet Patch DLin NLin AE Ocsvm IF PCA HBOS TFAD
CICIDS Aff-F 0.787 0.654 0.708 0.692 0.560 0.664 0.657 0.660 0.669 0.669 0.243 0.693 0.604 0.619 0.542 0.579
A-R 0.795 0.697 0.692 0.603 0.528 0.638 0.732 0.716 0.751 0.691 0.629 0.537 0.787 0.601 0.760 0.504
CalIt2 Aff-F 0.835 0.780 0.812 0.751 0.729 0.697 0.794 0.793 0.793 0.757 0.587 0.783 0.402 0.768 0.756 0.744
A-R 0.838 0.676 0.791 0.574 0.533 0.527 0.771 0.808 0.752 0.695 0.767 0.804 0.775 0.790 0.798 0.504
Credit Aff-F 0.750 0.744 0.713 0.663 0.650 0.632 0.744 0.746 0.738 0.742 0.561 0.714 0.634 0.710 0.695 0.600
A-R 0.958 0.957 0.934 0.703 0.552 0.504 0.957 0.957 0.954 0.948 0.909 0.953 0.860 0.871 0.951 0.500
GECCO Aff-F 0.908 0.893 0.839 0.701 0.782 0.687 0.894 0.906 0.893 0.882 0.823 0.666 0.424 0.785 0.708 0.627
A-R 0.970 0.952 0.795 0.714 0.516 0.555 0.954 0.949 0.947 0.936 0.769 0.804 0.619 0.711 0.557 0.499
Genesis Aff-F 0.896 0.833 0.891 0.810 0.856 0.776 0.864 0.856 0.856 0.829 0.854 0.677 0.788 0.814 0.721 0.535
A-R 0.974 0.676 0.690 0.937 0.947 0.659 0.913 0.685 0.696 0.755 0.931 0.733 0.549 0.815 0.897 0.497
MSL Aff-F 0.740 0.726 0.710 0.588 0.692 0.694 0.734 0.724 0.725 0.723 0.625 0.641 0.584 0.678 0.680 0.665
A-R 0.664 0.633 0.611 0.576 0.508 0.507 0.613 0.637 0.624 0.592 0.562 0.524 0.524 0.552 0.574 0.500
NYC Aff-F 0.994 0.769 0.684 0.708 0.853 0.862 0.794 0.776 0.828 0.819 0.689 0.667 0.648 0.680 0.675 0.689
A-R 0.816 0.466 0.640 0.633 0.671 0.549 0.791 0.709 0.768 0.671 0.504 0.456 0.475 0.666 0.446 0.502
PSM Aff-F 0.859 0.825 0.854 0.725 0.710 0.682 0.842 0.831 0.831 0.843 0.707 0.531 0.620 0.702 0.658 0.628
A-R 0.652 0.593 0.592 0.600 0.514 0.501 0.592 0.586 0.580 0.585 0.650 0.619 0.542 0.648 0.620 0.500
SMD Aff-F 0.847 0.840 0.827 0.679 0.724 0.675 0.831 0.845 0.841 0.844 0.439 0.742 0.626 0.738 0.629 0.660
A-R 0.811 0.722 0.745 0.631 0.508 0.502 0.727 0.736 0.728 0.738 0.774 0.602 0.664 0.679 0.626 0.500
ASD Aff-F 0.804 0.782 0.780 0.605 0.674 0.702 0.800 0.777 0.782 0.766 0.731 0.617 0.781 0.656 0.669 0.630
A-R 0.824 0.692 0.759 0.579 0.506 0.520 0.805 0.760 0.739 0.690 0.704 0.588 0.618 0.656 0.603 0.502
Contextual Aff-F 0.823 0.619 0.802 0.635 0.601 0.597 0.666 0.766 0.780 0.700 0.755 0.696 0.679 0.475 0.481 0.569
A-R 0.910 0.562 0.905 0.598 0.546 0.525 0.908 0.854 0.700 0.530 0.896 0.711 0.821 0.538 0.464 0.504
Global Aff-F 0.949 0.748 0.922 0.649 0.656 0.567 0.910 0.940 0.928 0.808 0.919 0.849 0.912 0.704 0.528 0.566
A-R 0.997 0.873 0.976 0.595 0.564 0.514 0.989 0.992 0.979 0.675 0.996 0.996 0.938 0.758 0.608 0.500
Seasonal Aff-F 0.997 0.681 0.992 0.776 0.788 0.859 0.992 0.989 0.993 0.951 0.927 0.805 0.938 0.637 0.673 0.686
A-R 0.998 0.512 0.946 0.701 0.584 0.644 0.958 0.922 0.823 0.623 0.949 0.829 0.918 0.437 0.516 0.502
Shapelet Aff-F 0.985 0.675 0.961 0.692 0.699 0.737 0.941 0.933 0.961 0.759 0.871 0.771 0.887 0.683 0.640 0.684
A-R 0.970 0.522 0.864 0.573 0.519 0.597 0.877 0.818 0.684 0.563 0.865 0.655 0.748 0.517 0.337 0.503
Trend Aff-F 0.916 0.734 0.901 0.677 0.584 0.765 0.897 0.888 0.721 0.830 0.699 0.691 0.914 0.693 0.669 0.642
A-R 0.892 0.612 0.847 0.524 0.500 0.569 0.858 0.835 0.671 0.642 0.482 0.471 0.878 0.484 0.468 0.502
Mixture Aff-F 0.892 0.856 0.862 0.652 0.641 0.709 0.863 0.879 0.727 0.839 0.673 0.676 0.881 0.676 0.667 0.710
A-R 0.931 0.763 0.854 0.570 0.522 0.516 0.861 0.863 0.767 0.749 0.493 0.475 0.911 0.517 0.531 0.501